# Towards Large Language Models with Self-Consistent Natural Language Explanations

### v2

@@ 3.4 Experimental Setup
We evaluate two instruction-tuned language models from the LLaMA family: LLaMA-3.1-8B-Instruct and LLaMA-3.2-3B-Instruct Touvron et al. (2023). All models are used in a zero-shot setting. To ensure compatibility with generation and attribution scoring, we add a padding token.
@@
We evaluate four diverse multiple-choice question-answering (MCQA) datasets: ECQA Aggarwal et al. (2021), ARC-Easy, ARC-Challenge Clark et al. (2018), and CODAH Chen et al. (2019). Each dataset contains 4–5 answer multiple-choice questions and is used to probe explanation consistency in different reasoning domains. We first preprocess each dataset into our unified attribution-enhanced format (Section 3.3) and then randomly split the resulting instances into the train (70%), validation (20%), and test (10%) sets using a fixed random seed. We use the training set to generate explanations and construct attribution-based preference pairs, the validation set for model selection, and the test set for final evaluation. For each training instance, we sample $k=5$ candidate explanations.
@@
To elicit decisions, we use a minimal task instruction prompt (e.g., “Choose the best plausible answer:”), and then generate post-hoc natural language explanations by conditioning on both the question $\mathbf{x}$ and the model’s answer $\mathbf{y}_{\mathrm{dec}}$. Explanations are generated using nucleus sampling with top-p = 0.9, temperature = 0.7, and a maximum output length of 400 tokens. Complete prompt templates are provided in Appendix A.
@@
We use the LIME feature attribution method Ribeiro et al. (2016), as implemented in Captum Kokhlikyan et al. (2020), to compute input token attributions with respect to the log-likelihood of the generated output sequence. We use 500 perturbation samples per example and set the reference input to the padding token repeated across the input length. Punctuation and formatting tokens (e.g., special symbols) are excluded from attribution using a manually defined skip list (see Appendix C).
@@
We compare the feature attribution vectors using two complementary alignment metrics that capture different aspects of self-consistency. Cosine Similarity quantifies how well an explanation’s attributions align with the decision’s direction and strength. Higher values indicate greater overlap in both the identity and importance of input features. Similar cosine-based comparisons were used in CC-SHAP Parcalabescu and Frank (2023). Spearman Rank Correlation assesses consistency in feature prioritization while abstracting away from attribution magnitudes:
$\mathrm{CC}_{\mathrm{sp}}=1-\frac{6\sum_{i=1}^{m}\left(\mathrm{r}\left(\phi^{(\mathbf{y}_{\mathrm{dec}})}_{i}\right)-\mathrm{r}\left(\phi^{(\mathbf{y}_{\mathrm{exp}})}_{i}\right)\right)^{2}}{m(m^{2}-1)}.$
Here, $\mathrm{r}(\cdot)$ assigns ordinal ranks to the attribution values, with rank 1 for the minimal score and $m$ for the largest; ties are resolved using average rank. This metric captures how similarly the two outputs prioritize input features, abstracting away from exact attribution values.
Together, these metrics capture different aspects of alignment: cosine similarity measures directional agreement, while Spearman captures feature ranking. High scores indicate that explanations highlight features similar to those of the decision, suggesting self-consistency; low scores imply divergent reasoning and reduced trust.
@@ 3.5 Key Findings and Insights
We evaluate attribution-based self-consistency across four MCQA datasets and two instruction-tuned language models, as shown in Table 1. We report decision accuracy for each model–dataset pair and compute self-consistency scores using five sampled explanations per input. Consistency is measured via cosine similarity (CC-LIME-Cos) and Spearman rank correlation (CC-LIME-Sp), reporting the worst, mean, and best alignment scores.
Explanation variability and metric differences.
We observe large gaps between the best and worst explanations per input. For instance, in ECQA with LLaMA3.1-8B, Spearman scores range from 5.07 to 31.76. Across settings, Spearman scores are generally higher than cosine scores, suggesting that models more consistently preserve feature ranking than exact attribution magnitudes. This greater variability makes Spearman particularly well-suited for identifying improvements, as it varies more significantly across explanations than cosine similarity, which tends to yield more compressed, lower values due to its sensitivity to attribution intensity. This variability confirms the need for a ranking-based evaluation and highlights the opportunity to identify highly self-consistent explanations from diverse generations (see Figure 3). Figure 4 illustrates that the large variance in Spearman correlation reflects meaningful semantic differences in explanation consistency.
Correctness and self-consistency.
To assess whether explanation self-consistency correlates with model correctness, we compare alignment scores for correct and incorrect predictions (Table 2). While some models show marginally higher self-consistency when correct, the differences are small and inconsistent. For instance, in ECQA with LLaMA3.1-8B, Spearman is slightly higher when correct (0.186 vs. 0.182), whereas in ARC-Easy with LLaMA3.2-3B, it is slightly higher when incorrect. Overall, we find no clear or robust trend linking correctness to explanation alignment, indicating that self-consistency captures a property orthogonal to prediction accuracy.
Model CC-Cos (C / I / $\Delta$) CC-Sp (C / I / $\Delta$)
@@
Table 2: Self-consistency scores (×100) for correct (C) and incorrect (I) predictions. We denote $\Delta=C-I.$ Cosine similarity and Spearman rank correlation are shown. Differences are small and vary in direction.
4 Improve Self-Consistency with DPO
Metric LLaMA3.1-8B LLaMA3.2-3B
[…]
ECQA CC-LIME-Cos Worst 07.58 ± 0.11 07.64 ± 0.12 ($\uparrow$ 0.8%) 01.11 ± 0.14 01.30 ± 0.13 ($\uparrow$ 17.1%)
Mean 08.05 ± 0.12 08.10 ± 0.12 ($\uparrow$ 0.6%) 01.46 ± 0.14 01.63 ± 0.13 ($\uparrow$ 11.6%)
Best 08.53 ± 0.13 08.57 ± 0.13 ($\uparrow$ 0.5%) 01.81 ± 0.14 01.96 ± 0.13 ($\uparrow$ 8.3%)
CC-LIME-Sp Worst 05.00 ± 0.30 07.20 ± 0.30 ($\uparrow$ 44%) 10.04 ± 0.31 10.50 ± 0.32 ($\uparrow$ 4.6%)
Mean 18.22 ± 0.23 20.22 ± 0.23 ($\uparrow$ 11%) 22.52 ± 0.26 22.65 ± 0.27 ($\uparrow$ 0.6%)
Best 31.30 ± 0.28 32.86 ± 0.29 ($\uparrow$ 5%) 35.02 ± 0.30 34.81 ± 0.30 ($\downarrow$ 0.6%)
ARC-Easy CC-LIME-Cos Worst 13.32 ± 0.25 13.47 ± 0.26 ($\uparrow$ 1.1%) 05.58 ± 0.32 05.58 ± 0.30
Mean 14.17 ± 0.27 14.31 ± 0.27 ($\uparrow$ 1.0%) 06.13 ± 0.33 06.11 ± 0.31 ($\downarrow$ 0.3%)
Best 15.03 ± 0.28 15.15 ± 0.29 ($\uparrow$ 0.8%) 06.68 ± 0.34 06.66 ± 0.32 ($\downarrow$ 2.9%)
CC-LIME-Sp Worst -01.90 ± 0.43 01.84 ± 0.44 ($\uparrow$ 96.8%) 07.97 ± 0.45 08.99 ± 0.44 ($\uparrow$ 12.8%)
Mean 12.01 ± 0.35 15.01 ± 0.38 ($\uparrow$ 25.0%) 20.72 ± 0.39 21.53 ± 0.39 ($\uparrow$ 3.9%)
Best 25.72 ± 0.43 28.38 ± 0.45 ($\uparrow$ 10.3%) 33.14 ± 0.44 34.19 ± 0.46 ($\uparrow$ 3.2%)
Table 3: Evaluation results across tasks. Other metrics show Worst, Best, and Mean self-consistency scores across 5 explanations (×100). Each model has a vanilla and a DPO-enhanced variant.
[…]
CODAH LLaMA3.1-8B None (Vanilla) 13.50 ± 0.28 19.46 ± 0.48
ECQA (DPO-tuned) 13.39 ± 0.27 ($\downarrow$ 0.8%) 21.58 ± 0.49 ($\uparrow$ 10.9%)
ARC-Easy (DPO-tuned) 13.38 ± 0.27 ($\downarrow$ 0.9%) 21.71 ± 0.50 ($\uparrow$ 11.6%)
LLaMA3.2-3B None (Vanilla) 08.04 ± 0.34 17.95 ± 0.57
ECQA (DPO-tuned) 08.16 ± 0.35 ($\uparrow$ 1.5%) 19.41 ± 0.54 ($\uparrow$ 8.1%)
ARC-Easy (DPO-tuned) 08.02 ± 0.34  ($\downarrow$ 0.2%) 18.13 ± 0.57 ($\uparrow$ 1.0%)
ARC-Chal LLaMA3.1-8B None (Vanilla) 16.26 ± 0.33 11.81 ± 0.45
ECQA (DPO-tuned) 16.48 ± 0.33 ($\uparrow$ 1.3%) 13.60 ± 0.49 ($\uparrow$ 13.16%)
ARC-Easy (DPO-tuned) 16.49 ± 0.33 ($\uparrow$ 1.4%) 13.87 ± 0.49 ($\uparrow$ 17.4%)
LLaMA3.2-3B None (Vanilla) 08.58 ± 0.46 19.51 ± 0.62
ECQA (DPO-tuned) 08.46 ± 0.44 ($\downarrow$ 1.4%) 19.83 ± 0.62 ($\uparrow$ 1.6%)
ARC-Easy (DPO-tuned) 08.49 ± 0.44 ($\downarrow$ 1.0%) 19.83 ± 0.69 ($\uparrow$ 1.6%)

### v3

@@ 3.4 Experimental Setup
We evaluate two instruction-tuned LLaMA models, LLaMA-3.1-8B-Instruct and LLaMA-3.2-3B-Instruct Touvron et al. (2023), in a zero-shot setting, using both the original Meta release and the Unsloth-optimized implementation (identical weights, minor decoding differences).
@@
Experiments cover four multiple-choice QA datasets: ECQA Aggarwal et al. (2021), ARC-Easy, ARC-Challenge Clark et al. (2018), and CODAH Chen et al. (2019), spanning diverse reasoning domains. Each dataset is converted into an attribution-enhanced format (Section 3.3) and split into train/validation/test (70%/20%/10%) with a fixed seed. Training instances are used to sample $k=5$ explanations and construct preference pairs, validation is used for model selection, and testing for final evaluation.
@@
Decisions are elicited using a minimal task instruction (e.g., “Choose the most plausible answer:”), and explanations are generated as in Section 3.3, conditioned on $(\mathbf{x},\mathbf{y}_{\mathrm{dec}})$. Explanations are sampled with nucleus decoding ($p=0.9$, $T=0.7$) and a maximum length of 400 tokens. The complete prompt templates are provided in Appendix A.
@@
We computed attributions using Local Interpretable Model-Agnostic Explanations (LIME) Ribeiro et al. (2016) and Layer Integrated Gradients (LIG) Sundararajan et al. (2017) as implemented in Captum Kokhlikyan et al. (2020), with sequence log-probability as the target. LIME used 500 perturbation samples with the padding token as baseline, and LIG used an embedding baseline with 25 interpolation steps. We applied the Unsloth variants for LIME computations and the original Meta models for LIG to ensure compatibility. Skip tokens were filtered as listed in Appendix C.
@@
We evaluated self-consistency with two alignment functions. Cosine similarity measures directional overlap between the attribution vectors, assessing whether explanations emphasize the same features with a similar magnitude. Spearman rank correlation captures agreement in feature prioritization while abstracting from the attribution scale:
$\mathrm{CC}_{\mathrm{sp}}=1-\frac{6\sum_{i=1}^{m}\left(\mathrm{r}\!\left(\phi^{(\mathbf{y}_{\mathrm{dec}})}_{i}\right)-\mathrm{r}\!\left(\phi^{(\mathbf{y}_{\mathrm{exp}})}_{i}\right)\right)^{2}}{m(m^{2}-1)},$
where $\mathrm{r}(\cdot)$ assigns the ordinal ranks (ties are decided by average). High values indicate that explanations highlight features aligned with the decision.
@@ 3.5 Key Findings and Insights
We report self-consistency across all model–dataset pairs using both LIME and LIG attributions, measured by cosine similarity (CC-Cos) and Spearman rank correlation (CC-Sp). For each setting, we report the worst, mean, and best scores across five sampled explanations.
LLaMA3.1-8B LLaMA3.2-3B
LIME LIG LIME LIG
ECQA (#10882) Acc. 71.11 66.51 65.85 64.83
$C$-Cos Worst 07.70 01.36 01.30 01.82
Mean 08.18 01.67 01.65 02.25
Best 08.67 01.98 01.99 02.66
$C$-Sp Worst 05.07 33.79 09.75 30.55
Mean 18.47 41.70 22.28 38.72
Best 31.76 49.11 34.64 46.54
ARC-E (#5197) Acc. 87.01 82.87 81.90 80.69
$C$-Cos Worst 13.31 01.49 05.32 01.65
Mean 14.16 01.78 05.86 02.04
Best 15.02 02.07 06.41 02.41
$C$-Sp Worst −01.06 38.81 07.26 36.63
Mean 12.44 46.42 19.68 45.58
Best 25.77 53.41 31.98 53.68
ARC-C (#2590) Acc. 77.57 71.12 68.70 64.90
$C$-Cos Worst 15.09 01.39 07.67 01.49
Mean 16.07 01.67 08.29 01.90
Best 17.08 01.94 08.93 02.29
$C$-Sp Worst −01.82 37.07 05.94 33.22
Mean 11.48 44.87 18.44 41.43
Best 24.73 52.11 30.89 49.12
CODAH (#2776) Acc. 83.39 75.48 72.05
$C$-Cos Worst 12.87 07.49 01.91
Mean 13.68 08.06 02.47
Best 14.52 08.64 02.96
$C$-Sp Worst 07.11 06.82 35.24
Mean 19.97 19.39 44.88
Best 32.66 31.95 53.60
Table 1: PSCB self-consistency scores ($\times 100$) on the test split of each dataset, for both LLaMA models and both attribution methods (LIME, LIG). Worst, Mean, and Best are reported over 5 sampled explanations per item. $C$-Cos: cosine similarity; $C$-Sp: Spearman rank correlation between decision and explanation attribution vectors.
Model CC-Cos (T / F / $\Delta$) CC-Sp (T / F / $\Delta$)
@@
Table 2: Self-consistency scores (×100) for true (T) and false (F) predictions. We denote $\Delta=T-F.$ Cosine similarity and Spearman rank correlation are shown. Differences are small and vary in direction.
Explanation variability and metric differences.
We observe substantial gaps between the best and worst explanations for a given input. For example, in ECQA with LLaMA-3.1-8B (LIME attributions), Spearman scores range from 5.07 to 31.76. Across datasets and models, Spearman correlations are consistently higher than cosine similarities, indicating that models preserve feature ranking more reliably than attribution magnitudes. Spearman also exhibits greater spread across explanations, making it more effective in detecting consistency differences, whereas cosine tends to produce compressed values due to its sensitivity to attribution scale. This supports the need for ranking-based evaluation and highlights the opportunity to identify highly self-consistent explanations among diverse generations (Figure 3). Figure 4 further shows that a large Spearman variance corresponds to meaningful semantic differences in explanation quality.
Correctness and self-consistency.
We test whether self-consistency correlates with prediction correctness by comparing alignment scores for correct vs. incorrect predictions (Table 2 in Appendix). Differences are small and inconsistent across datasets and models, suggesting that self-consistency is largely orthogonal to accuracy.
4 Improving Self-Consistency with DPO
  LLaMA3.1-8B LLaMA3.2-3B
  B-B B-T ($\uparrow\downarrow$) T-T ($\uparrow\downarrow$)   B-B B-T ($\uparrow\downarrow$) T-T ($\uparrow\downarrow$)
ECQA LIME Acc.    70.34 70.34 69.70 $\downarrow$0.9%   67.13 67.13 66.12 $\downarrow$1.5%
$C$-Cos Worst    07.58 $\pm 0.11$ 07.64 $\pm 0.12$ $\uparrow$0.8% 09.16 $\pm 0.12$ $\uparrow$20.8%   01.11 $\pm 0.14$ 01.30 $\pm 0.13$ $\uparrow$17.1% 02.14 $\pm 0.13$ $\uparrow$92.8%
Mean    08.05 $\pm 0.12$ 08.10 $\pm 0.12$ $\uparrow$0.6% 09.72 $\pm 0.12$ $\uparrow$20.7%   01.46 $\pm 0.14$ 01.63 $\pm 0.13$ $\uparrow$11.6% 02.47 $\pm 0.14$ $\uparrow$69.2%
Best    08.53 $\pm 0.13$ 08.57 $\pm 0.13$ $\uparrow$0.5% 10.28 $\pm 0.13$ $\uparrow$20.5%   01.81 $\pm 0.14$ 01.96 $\pm 0.13$ $\uparrow$8.3% 02.80 $\pm 0.14$ $\uparrow$54.7%
$C$-Sp Worst    05.00 $\pm 0.30$ 07.20 $\pm 0.30$ $\uparrow$44.0% 07.86 $\pm 0.29$ $\uparrow$57.2%   10.04 $\pm 0.31$ 10.50 $\pm 0.32$ $\uparrow$4.6% 12.94 $\pm 0.32$ $\uparrow$28.9%
Mean    18.22 $\pm 0.23$ 20.22 $\pm 0.23$ $\uparrow$11.0% 20.58 $\pm 0.24$ $\uparrow$13.0%   22.52 $\pm 0.26$ 22.65 $\pm 0.27$ $\uparrow$0.6% 24.71 $\pm 0.27$ $\uparrow$9.7%
Best    31.30 $\pm 0.28$ 32.86 $\pm 0.29$ $\uparrow$5.0% 33.45 $\pm 0.28$ $\uparrow$6.9%   35.02 $\pm 0.30$ 34.81 $\pm 0.30$ $\downarrow$0.6% 36.50 $\pm 0.29$ $\uparrow$4.2%
LIG Acc.    67.03 67.03 67.68 $\uparrow$1.1%   62.44 62.44 62.99 $\uparrow$0.9%
$C$-Cos Worst    0.87 $\pm$ 0.03 01.35 $\pm$ 0.02 $\uparrow$55.2% 01.36 $\pm$ 0.02 $\uparrow$56.3%   01.90 $\pm$ 0.03 01.97 $\pm$ 0.03 $\uparrow$3.7% 02.54 $\pm$ 0.03 $\uparrow$33.7%
Mean    01.22 $\pm$ 0.03 01.67 $\pm$ 0.02 $\uparrow$36.9% 01.67 $\pm$ 0.02 $\uparrow$36.9%   02.32 $\pm$ 0.03 02.37 $\pm$ 0.03 $\uparrow$2.2% 02.82 $\pm$ 0.03 $\uparrow$21.6%
Best    01.55 $\pm$ 0.03 01.96 $\pm$ 0.03 $\uparrow$26.5% 01.96 $\pm$ 0.02 $\uparrow$26.5%   02.72 $\pm$ 0.03 02.73 $\pm$ 0.04 $\uparrow$0.4% 02.99 $\pm$ 0.03 $\uparrow$9.9%
$C$-Sp Worst    34.71 $\pm$ 0.44 34.71 $\pm$ 0.45 36.59 $\pm$ 0.42 $\uparrow$5.4%   28.61 $\pm$ 0.45 29.04 $\pm$ 0.45 $\uparrow$1.5% 33.30 $\pm$ 0.45 $\uparrow$16.4%
Mean    42.52 $\pm$ 0.40 42.65 $\pm$ 0.41 $\uparrow$0.3% 43.66 $\pm$ 0.39 $\uparrow$2.7%   36.72 $\pm$ 0.42 37.17 $\pm$ 0.41 $\uparrow$1.2% 41.90 $\pm$ 0.41 $\uparrow$14.1%
Best    49.87 $\pm$ 0.40 50.15 $\pm$ 0.40 $\uparrow$0.6% 51.24 $\pm$ 0.38 $\uparrow$2.7%   44.53 $\pm$ 0.41 45.05 $\pm$ 0.41 $\uparrow$1.2% 50.17 $\pm$ 0.41 $\uparrow$12.7%
ARC-Easy LIME Acc.    87.72 87.72 86.76 $\downarrow$1.1%   81.62 81.62 81.04 $\downarrow$0.7%
$C$-Cos Worst    13.32 $\pm 0.25$ 13.47 $\pm 0.26$ $\uparrow$1.1% 14.97 $\pm 0.24$ $\uparrow$12.4%   05.58 $\pm 0.32$ 05.58 $\pm 0.30$ 06.32 $\pm 0.29$ $\uparrow$13.2%
Mean    14.17 $\pm 0.27$ 14.31 $\pm 0.27$ $\uparrow$1.0% 15.89 $\pm 0.25$ $\uparrow$12.1%   06.13 $\pm 0.33$ 06.11 $\pm 0.31$ $\downarrow$0.3% 06.87 $\pm 0.30$ $\uparrow$12.1%
Best    15.03 $\pm 0.28$ 15.15 $\pm 0.29$ $\uparrow$0.8% 16.82 $\pm 0.27$ $\uparrow$11.9%   06.68 $\pm 0.34$ 06.66 $\pm 0.32$ $\downarrow$2.9% 07.42 $\pm 0.31$ $\uparrow$11.1%
$C$-Sp Worst    -01.90 $\pm 0.43$ 01.84 $\pm 0.44$ $\uparrow$96.8% 01.76 $\pm 0.44$ $\uparrow$92.6%   07.97 $\pm 0.45$ 08.99 $\pm 0.44$ $\uparrow$12.8% 11.48 $\pm 0.46$ $\uparrow$44.0%
Mean    12.01 $\pm 0.35$ 15.01 $\pm 0.38$ $\uparrow$25.0% 14.88 $\pm 0.37$ $\uparrow$23.9%   20.72 $\pm 0.39$ 21.53 $\pm 0.39$ $\uparrow$3.9% 23.16 $\pm 0.39$ $\uparrow$11.8%
Best    25.72 $\pm 0.43$ 28.38 $\pm 0.45$ $\uparrow$10.3% 27.86 $\pm 0.44$ $\uparrow$8.3%   33.14 $\pm 0.44$ 34.19 $\pm 0.46$ $\uparrow$3.2% 34.80 $\pm 0.41$ $\uparrow$5.0%
LIG Acc.    84.26 84.26 84.84 $\uparrow$0.7%   80.23 80.23 79.27 $\downarrow$1.2%
$C$-Cos Worst    0.98 $\pm$ 0.05 01.56 $\pm$ 0.04 $\uparrow$59.2% 01.54 $\pm$ 0.04 $\uparrow$57.1%   01.20 $\pm$ 0.04 01.25 $\pm$ 0.04 $\uparrow$4.2% 01.71 $\pm$ 0.04 $\uparrow$42.5%
Mean    01.35 $\pm$ 0.05 01.84 $\pm$ 0.04 $\uparrow$36.3% 01.82 $\pm$ 0.04 $\uparrow$34.8%   01.62 $\pm$ 0.04 01.66 $\pm$ 0.04 $\uparrow$2.5% 02.07 $\pm$ 0.04 $\uparrow$27.8%
Best    01.68 $\pm$ 0.05 02.11 $\pm$ 0.04 $\uparrow$25.6% 02.08 $\pm$ 0.04 $\uparrow$23.8%   02.01 $\pm$ 0.04 02.03 $\pm$ 0.04 $\uparrow$1.0% 02.43 $\pm$ 0.04 $\uparrow$20.9%
$C$-Sp Worst    37.40 $\pm$ 0.64 38.94 $\pm$ 0.66 $\uparrow$4.1% 39.93 $\pm$ 0.65 $\uparrow$6.8%   33.77 $\pm$ 0.65 34.88 $\pm$ 0.67 $\uparrow$3.3% 37.97 $\pm$ 0.65 $\uparrow$12.4%
Mean    45.55 $\pm$ 0.55 46.76 $\pm$ 0.57 $\uparrow$2.7% 47.09 $\pm$ 0.58 $\uparrow$3.4%   43.15 $\pm$ 0.60 44.02 $\pm$ 0.61 $\uparrow$2.0% 46.24 $\pm$ 0.60 $\uparrow$7.2%
Best    52.88 $\pm$ 0.52 53.63 $\pm$ 0.53 $\uparrow$1.4% 53.79 $\pm$ 0.55 $\uparrow$1.7%   51.75 $\pm$ 0.60 52.52 $\pm$ 0.60 $\uparrow$1.5% 53.74 $\pm$ 0.60 $\uparrow$3.8%
Table 3: In-Domain Performance: Models trained and evaluated on the same dataset (ECQA, ARC-Easy). Results are based on LIME and LIG attributions, reporting Worst, Mean, and Best self-consistency scores across 5 sampled explanations (×100). Accuracy is shown alongside. Each model is evaluated under three modes (B–B, B–T, T–T), with relative improvements over B–B indicated by arrows.
Training Data  Acc.   $C$-Cos ($\uparrow\downarrow$) $C$-Sp ($\uparrow\downarrow$)
    B-T T-T    B-T T-T
CODAH L3.1-8B None    81.29   13.50 ± 0.28    19.46 ± 0.48
ECQA    77.70   13.39 ± 0.27 ($\downarrow$ 0.8%) 14.05 ± 0.32 ($\uparrow$ 4.1%)    18.30 ± 0.56 ($\downarrow$ 6.0%) 21.58 ± 0.49 ($\uparrow$ 10.9%)
ARC-Easy    80.94   13.38 ± 0.27 ($\downarrow$ 0.8%) 14.29 ± 0.29 ($\uparrow$ 5.9%)    18.79 ± 0.57 ($\downarrow$ 3.4%) 21.71 ± 0.50 ($\uparrow$ 11.6%)
L3.2-3B None    73.90   08.04 ± 0.34    17.95 ± 0.57
ECQA    72.43   08.16 ± 0.35 ($\uparrow$ 1.5%) 09.42 ± 0.31 ($\uparrow$ 17.2%)    19.41 ± 0.54 ($\uparrow$ 8.1%) 20.10 ± 0.55 ($\uparrow$ 12.0%)
ARC-Easy    74.63   08.02 ± 0.34  ($\downarrow$ 0.2%) 09.51 ± 0.31 ($\uparrow$ 18.3%)    18.13 ± 0.57 ($\uparrow$ 1.0%) 20.52 ± 0.55 ($\uparrow$ 14.3%)
ARC-Chal L3.1-8B None    76.15   16.26 ± 0.33    11.81 ± 0.45
ECQA    73.85   16.48 ± 0.33 ($\uparrow$ 1.3%) 17.57 ± 0.32 ($\uparrow$ 8.1%)    13.60 ± 0.49 ($\uparrow$ 13.16%) 13.45 ± 0.49 ($\uparrow$ 13.9%)
ARC-Easy    73.46   16.49 ± 0.33 ($\uparrow$ 1.4%) 17.65 ± 0.31 ($\uparrow$ 8.5%)    13.87 ± 0.49 ($\uparrow$ 17.4%) 13.31 ± 0.49 ($\uparrow$ 12.7%)
L3.2-3B None    66.67   08.58 ± 0.46    19.51 ± 0.62
ECQA    66.67   08.46 ± 0.44 ($\downarrow$ 1.4%) 09.13 ± 0.41 ($\uparrow$ 6.4%)    19.83 ± 0.62 ($\uparrow$ 1.6%) 21.37 ± 0.59 ($\uparrow$ 9.5%)
ARC-Easy    67.06   08.49 ± 0.44 ($\downarrow$ 1.0%) 09.08 ± 0.42 ($\uparrow$ 5.8%)    19.83 ± 0.69 ($\uparrow$ 1.6%) 21.78 ± 0.59 ($\uparrow$ 11.6%)